# Do Students Write Better Post-AI Support?

### v1

@@ 3.2 Participants
A total of 81 participants completed the study. However, two participants were excluded from the analysis due to providing written responses shorter than 20 words, resulting in a final sample of 79 participants. This exceeded the minimum sample size (72 participants; 36 per condition) required by a priori power analysis (effect size , power = 0.80, ). Participants were recruited via Prolific111https://www.prolific.co/ and compensated £8 for approximately one hour of participation. Participants were randomly assigned to one of two experimental conditions: passive chatbot (39) or proactive chatbot (40). All participants were current students or graduates from medical or nursing programs. The sample comprised 45 females and 34 males, primarily aged 18-24 (28) and 25-34 (31); fewer participants belonged to age groups 35-44 (11), 45-54 (6), 55-64 (2), and 65 or older (1). Participants were distributed geographically across North/Central America (34), Europe (21), Africa (15), Australia (4), South America (2), and other regions (3). Educational backgrounds included Bachelor’s degrees (36), high school diploma or equivalent (18), Master’s degrees (11), vocational training or diplomas (7), Doctoral degrees (4), and other qualifications (3). Familiarity with GenAI tools was predominantly intermediate (42), beginner (25), advanced (9), expert (2), and none (1).
@@ 3.3.1 Learning Materials
The multimodal data were used to generate three visuals (Figure 1), each with progressively increasing complexity. First, a bar chart summarises time spent on different behaviours derived from positional data. This visual information illustrates team-level resource allocation and prioritisation strategies 64. Second, a communication network diagram combining positional and audio data to reveal the frequency and direction of verbal interactions among nurses, the doctor, the patient, and the patient’s relative, thereby depicting participant roles and their level of involvement 68. Third, a ward map integrates positional, audio, and physiological data. Inspired by visual and spatial analytics methods used in professional sports 20, the ward map utilises heatmap visualisation techniques to show communication intensity, trace spatial movements, and highlight locations of peak physiological arousal via heart rate data.
@@ 3.3.2 Research Platform
The academic writing tasks were conducted using a custom-built platform comprising three main components. The first component, the display area (Figure 2), presented one visual analytic at a time to minimise visual clutter and prevent information overload 14. Participants could freely navigate between the three visual information representations using the navigation arrows positioned on either side of the component. The second component, the writing space (Figure 2), provided participants with an area to articulate their analytical insights by synthesising information presented in the visual information. The third component varied based on the experimental condition. In the AI-supported conditions, participants had access to a chat function (Figure 2), offering real-time support from GenAI agents, which responded either passively or proactively. Conversely, in conditions without AI support, the chat function was replaced by an instruction component providing standard task instruction without AI assistance. The overall platform design emphasised simplicity 52, enabling participants to focus primarily on comprehending the visual information and effectively completing the tasks without being overwhelmed by complex functionalities.
@@ 3.3.3 Chatbot Interaction Design
[…]
Building upon the VizChat prototype 65, both chatbots adopted an advanced agentic architecture composed of four interconnected modules. Prompts were categorised into passive and proactive types corresponding to the aforementioned interaction strategies. At its core, the GenAI agent managed and synthesised inputs from visual information and user prompts, utilising Retrieval-Augmented Generation (RAG) to efficiently retrieve data-informed content from its knowledge database 19. This comprehensive knowledge repository contained contextual information, visualisation details, and interaction logs that enabled personalised and relevant interactions. Finally, synthesised prompts were processed by the multimodal GPT-4o AI model, generating coherent and context-aware responses that enriched the database and progressively improved the quality of subsequent learner interactions.
@@ 3.4.1 Generative AI Measurement (GLAT)
We assessed learners’ generative AI literacy using the Generative AI Literacy Assessment Test (GLAT) 28. This 20-item instrument was selected for its established validity and reliability. Face validity was ensured through expert review and revision, while content validity was evaluated with 204 participants to confirm item relevance, comprehensiveness and comprehensibility. Structural validity and internal consistency were confirmed using classical test theory and item response theory, supporting a reliable two-parameter logistic (2PL) model (Cronbach’s ) and robust factor structure (, , , 90% CI ). As the 2PL model accounts only for item difficulty and discrimination, we applied a correction-for-guessing formula 54 to control for potential guessing effects (Eq. 1).
In this formula, represents the final score corrected for guessing behaviours, denotes the number of correctly answered items by a given participant, denotes the number of wrongly answered items, and denotes the number of answer choices per item, which is four for the GLAT.
@@ 3.4.2 Domain Knowledge Measurement (VLAT)
Visualisation literacy constitutes essential domain knowledge for effective multimodal writing 17, 37. We assessed participants’ domain knowledge in visualisation literacy using the Visualisation Literacy Assessment Test (mini-VLAT) 48. This 12-item instrument was selected for its established validity and reliability, which are comparable to those of the original 53-item VLAT 37, while providing a briefer assessment suitable for our extended study. Its concise structure helped minimise participants’ cognitive load. Structural validity and internal consistency for this instrument have previously been confirmed 48. To account for potential guessing effects, we also applied the correction-for-guessing formula 54 to participants’ scores.
@@ 3.4.3 Academic Writing Assessment Rubric
A comprehensive assessment rubric was developed to evaluate the participants’ academic writing performances in tasks involving visual data analyses. Based on prior literature and established theoretical grounding within academic writing research (as elaborated in Section 2.1), we identified five distinct evaluation metrics (Table 1): Insightfulness,Visual Data Integration, Organisation and Coherence, Linguistic Quality, and Critical Thinking. Each criterion was evaluated using a 5-point Likert scale (Level 1–Poor to Level 5–Excellent), with comprehensive descriptors and examples available in the full rubric (see Supplementary Information).
@@
| Insightfulness | Assesses the originality, depth, and relevance of insights. High-quality responses identify subtle patterns, offer holistic interpretations, and provide conclusions that extend beyond obvious points. Low-quality responses are surface-level observations or simplistic interpretations 32, 9. | 0.82 |
| Visual Data Integration | Measures the effectiveness and clarity with which visual data (charts, graphs, communication networks, heart rate information) is explicitly interpreted, combined, and connected into the written analysis. High-quality responses synthesize multiple visual data sources cohesively, while low-quality responses merely mention data without synthesizing or miss integrating data entirely 27. | 0.85 |
| Organisation and Coherence | Evaluates the clarity, logical structure, and connectedness of ideas presented in the writing. High-quality responses have logical structure, smooth transitions between points, and clearly organized arguments. Low-quality responses lack clarity, exhibit abrupt transitions, or are disorganized 47. | 0.75 |
| Linguistic Quality | Reflects grammar accuracy, vocabulary appropriateness, sentence fluency, and overall clarity of written expression. High-quality responses utilize clear, precise, and varied language with minimal errors, while low-quality responses contain grammatical errors, limited vocabulary, or unclear expression 5. | 0.89 |
| Critical Thinking | Evaluates the level of careful thought, complexity, consideration of alternative viewpoints, and strength of conclusions and recommendations. High-quality responses deeply analyze complexities and implications, considering broader contexts, contingencies, or solutions. Low-quality responses show limited or superficial analytical engagement and fail to address alternative perspectives or implications 15, 43. | 0.80 |
@@ 3.4.4 Scoring Procedures and Inter-Rater Reliability
The scoring process followed detailed guidelines based on the rubric provided in the Supplementary Information. Two independent, trained coders rated each participant’s writing samples at each study phase (AI-support and AI-removal). Coders were trained extensively on the rubric definitions through practice scoring sessions with sample responses (n=10), followed by discussion and calibration of understanding to maximise rating consistency. All essays were then independently blind-rated by both coders. After ratings were fully completed, reliability analyses were performed (Spearman’s correlations), demonstrating high inter-rater reliability ( in all cases; Table 1). Final participant scores for each dimension, at each stage, were averaged across the two coders, producing a single composite rating for each dimension and subsequently an overall Composite Score. To ensure transparency and replicability, example excerpts representing each scoring level were provided in the detailed rubric (see Supplementary Information).
@@ 3.5 Study Procedure
We conducted a two-phase, mixed-method experimental study to examine how GenAI literacy and AI interaction type (passive vs. proactive chatbot) influenced participants’ academic writing performance with visual information. Participants were recruited via Prolific and directed to a Qualtrics222https://www.qualtrics.com/ survey, where they provided informed consent, demographics, and completed literacy assessments (GLAT and mini-VLAT). After reviewing contextual information on the healthcare simulation (Section 3.3.1), they were randomly assigned to either the passive or proactive chatbot assistance condition. In the AI-supported phase, participants composed a 150-word written response interpreting the provided multimodal visuals with real-time chatbot support. Subsequently, in the AI-removal phase, participants completed a similar task with new visual data from a different simulation scenario without chatbot access (Figure 4). The study lasted approximately one hour. Ethical approval was obtained from Monash University (Project Number: 37307).
@@ 3.6.1 Correlation Analysis (RQ1)
To examine the relationship between participants’ GenAI literacy and academic writing performance (RQ1), correlation analyses were conducted separately for each chatbot engagement condition (passive, proactive), at both AI-support and AI-removal time points, and across the five writing dimensions, as well as the composite scores. GenAI literacy was first screened for normality to determine the suitable statistical test. The Shapiro-Wilk normality tests indicated a violation of normality assumptions for GenAI literacy (). Thus, Spearman’s rank-order correlation was used for all analyses. For each condition separately, Spearman’s correlations were calculated between participants’ GenAI literacy scores and their average writing scores at the AI-support and AI-removal stages.
@@ 3.6.2 Comparative Analysis (RQ2)
To investigate further if students at varying levels of GenAI literacy performed differently in academic writing (RQ2), we conducted comparative analyses by grouping participants into two distinct GenAI literacy categories: low and high literacy. Participants were categorised based on median splits of GenAI literacy scores within each condition separately. We chose the median split procedure to ensure balanced group sizes and improve interoperability. To ascertain whether academic writing performance differed significantly between low and high GenAI literacy groups, we used Mann-Whitney U tests due to the ordinal nature of the rating scales and violations of normality assumptions (verified through Shapiro-Wilk tests). Analyses were conducted separately for each evaluation criterion (insightfulness, visual data integration, organisation and coherence, linguistic quality, critical thinking) as well as for the composite score, for both the AI-support and AI-removal stages periods. Statistical significance was set at and adjusted for multiple comparisons using the Bonferroni method. Significant differences were further interpreted using effect sizes calculated as the Rank-Biserial correlation.
@@ 3.6.3 Regression Analysis (RQ3)
To investigate the extent GenAI literacy predicted writing performance after controlling for domain-specific knowledge (RQ3), ordinal logistic regression analyses were conducted. As shown in the following formula (1), the dependent variables for regression analyses were participants’ academic writing performance scores during the AI-removal period for each dimension (insightfulness, visual data integration, organisation and coherence, linguistic quality, critical thinking) separately, as well as the composite writing score. Each outcome variable, measured using ordinal scales (; scores from 1 to 5; refers to each ordinal threshold, ), justified the use of ordinal logistic regression rather than linear regression. Primary predictor variables included standardised GenAI literacy scores (), domain knowledge (; VLAT score), chatbot intervention type (; binary-coded: 0 = passive chatbot, 1 = proactive chatbot), and participants’ AI-assisted scores (; to control writing status during chatbot support availability). Additionally, denotes the threshold/intercept parameter for each category boundary, and , and represent regression coefficients for the predictors.

### v2

@@ 3.2 Participants
A total of 81 participants completed the study. However, two participants were excluded from the analysis due to providing written responses shorter than 20 words, resulting in a final sample of 79 participants. This exceeded the minimum sample size (72 participants; 36 per condition) required by a priori power analysis (effect size , power = 0.80, ). Participants were recruited via Prolific111https://www.prolific.co/ and compensated £8 for approximately one hour of participation. Participants were randomly assigned to one of two experimental conditions: reactive chatbot (39) or proactive chatbot (40). All participants were current students or graduates from medical or nursing programs. The sample comprised 45 females and 34 males, primarily aged 18-24 (28) and 25-34 (31); fewer participants belonged to age groups 35-44 (11), 45-54 (6), 55-64 (2), and 65 or older (1). Participants were distributed geographically across North/Central America (34), Europe (21), Africa (15), Australia (4), South America (2), and other regions (3). Educational backgrounds included Bachelor’s degrees (36), high school diploma or equivalent (18), Master’s degrees (11), vocational training or diplomas (7), Doctoral degrees (4), and other qualifications (3). Familiarity with GenAI tools was predominantly intermediate (42), beginner (25), advanced (9), expert (2), and none (1).
@@ 3.3.1 Learning Materials
The multimodal data were used to generate three visuals (Figure 1), each with progressively increasing complexity. First, a bar chart summarises time spent on different behaviours derived from positional data. This visual information illustrates team-level resource allocation and prioritisation strategies [Anonymised]. Second, a communication network diagram combining positional and audio data to reveal the frequency and direction of verbal interactions among nurses, the doctor, the patient, and the patient’s relative, thereby depicting participant roles and their level of involvement [Anonymised]. Third, a ward map integrates positional, audio, and physiological data. Inspired by visual and spatial analytics methods used in professional sports (Goldsberry, 2012), the ward map utilises heatmap visualisation techniques to show communication intensity, trace spatial movements, and highlight locations of peak physiological arousal via heart rate data.
@@ 3.3.2 Research Platform
The academic writing tasks were conducted using a custom-built platform comprising three main components. The first component, the display area (Figure 2), presented one visual analytic at a time to minimise visual clutter and prevent information overload (Ellis and Dix, 2007). Participants could freely navigate between the three visual information representations using the navigation arrows positioned on either side of the component. The second component, the writing space (Figure 2), provided participants with an area to articulate their analytical insights by synthesising information presented in the visual information. The third component varied based on the experimental condition. In the AI-supported conditions, participants had access to a chat function (Figure 2), offering real-time support from GenAI agents, which responded either reactively or proactively. Conversely, in conditions without AI support, the chat function was replaced by an instruction component providing standard task instruction without AI assistance. The overall platform design emphasised simplicity (Stone et al., 2005), enabling participants to focus primarily on comprehending the visual information and effectively completing the tasks without being overwhelmed by complex functionalities.
@@ 3.3.3 Chatbot Interaction Design
We developed two distinct GenAI chatbots, namely the reactive and the proactive, primarily differing in their interaction types and learner-support strategies (Figure 3). The purpose behind these design choices was to explore how interaction type associates with the writing performance of learners with varying GenAI literacy levels. While both chatbots introduced users with a welcome message explaining the writing task, the reactive chatbot provided a description of its reactive and supportive features and responds only to users’ questions without following up with further interactions (Ma et al., 2023; Yan et al., 2024b). In contrast, the proactive chatbot actively initiated conversations with users and continually engaged them through domain expert-designed questions and structured narrative pathways (see Appendix D) (Park et al., 2023). Rather than simply reacting, the proactive chatbot offered structured questioning sequences and iterative feedback, progressively enhancing learners’ understanding of visual information based on Hattie and Timperley (2007)’s feedback model (e.g., task and process-level). For example, when guiding learners through the ward map analyses, the proactive chatbot first prompted them towards interpreting the distribution of hexagons in the ward (positional data). Once it confirmed that the user had clearly understood this aspect, it then moved onto the next layer of interpretation, specifically the intensity of the colour (verbal communication). This scaffolded questioning approach aligned closely with established conceptual guidelines for interpreting visual information (Xun and Land, 2004), systematically aiming to enhance learners’ critical interpretation skills.
Building upon the VizChat prototype (Yan et al., 2024b), both chatbots adopted an advanced agentic architecture composed of four interconnected modules. Prompts were categorised into reactive and proactive types corresponding to the aforementioned interaction strategies. At its core, the GenAI agent managed and synthesised inputs from visual information and user prompts, utilising Retrieval-Augmented Generation (RAG) to efficiently retrieve data-informed content from its knowledge database (Gao et al., 2023). This comprehensive knowledge repository contained contextual information, visualisation details, and interaction logs that enabled personalised and relevant interactions. Finally, synthesised prompts were processed by the multimodal GPT-4o AI model, generating coherent and context-aware responses that enriched the database and progressively improved the quality of subsequent learner interactions.
@@ 3.4.1 Generative AI Measurement (GLAT)
We assessed learners’ generative AI literacy using the Generative AI Literacy Assessment Test (GLAT) (Jin et al., 2024). This 20-item instrument was selected for its established validity and reliability. Face validity was ensured through expert review and revision, while content validity was evaluated with 204 participants to confirm item relevance, comprehensiveness and comprehensibility. Structural validity and internal consistency were confirmed using classical test theory and item response theory, supporting a reliable two-parameter logistic (2PL) model (Cronbach’s ) and robust factor structure (, , , 90% CI ). As the 2PL model accounts only for item difficulty and discrimination, we applied a correction-for-guessing formula (Thorndike et al., 1991) to control for potential guessing effects (Eq. 1).
In this formula, represents the final score corrected for guessing behaviours, denotes the number of correctly answered items by a given participant, denotes the number of wrongly answered items, and denotes the number of answer choices per item, which is four for the GLAT. The full test is available in Appendix A.
@@ 3.4.2 Domain Knowledge Measurement (VLAT)
Visualisation literacy constitutes essential domain knowledge for effective multimodal writing (Felten, 2008; Lee et al., 2016). We assessed participants’ domain knowledge in visualisation literacy using the Visualisation Literacy Assessment Test (mini-VLAT) (Pandey and Ottley, 2023). This 12-item instrument was selected for its established validity and reliability, which are comparable to those of the original 53-item VLAT (Lee et al., 2016), while providing a briefer assessment suitable for our extended study. Its concise structure helped minimise participants’ cognitive load. Structural validity and internal consistency for this instrument have previously been confirmed (Pandey and Ottley, 2023). To account for potential guessing effects, we also applied the correction-for-guessing formula (Thorndike et al., 1991) to participants’ scores. The types of visualisations and their corresponding questions are provided in Appendix B.
@@ 3.4.3 Academic Writing Assessment Rubric
A comprehensive assessment rubric was developed to evaluate the participants’ academic writing performances in tasks involving visual data analyses. Based on prior literature and established theoretical grounding within academic writing research (as elaborated in Section 2.1), we identified five distinct evaluation metrics (Table 1): Insightfulness,Visual Data Integration, Organisation and Coherence, Linguistic Quality, and Critical Thinking. Each criterion was evaluated using a 5-point Likert scale (Level 1–Poor to Level 5–Excellent), with comprehensive descriptors and examples available in the full rubric (see Appendix C).
@@
| Insightfulness | Assesses the originality, depth, and relevance of insights. High-quality responses identify subtle patterns, offer holistic interpretations, and provide conclusions that extend beyond obvious points. Low-quality responses are surface-level observations or simplistic interpretations (Kuhn, 2005; Chanock, 2000). | 0.82 |
| Visual Data Integration | Measures the effectiveness and clarity with which visual data (charts, graphs, communication networks, heart rate information) is explicitly interpreted, combined, and connected into the written analysis. High-quality responses synthesize multiple visual data sources cohesively, while low-quality responses merely mention data without synthesizing or miss integrating data entirely (Jewitt, 2006). | 0.85 |
| Organisation and Coherence | Evaluates the clarity, logical structure, and connectedness of ideas presented in the writing. High-quality responses have logical structure, smooth transitions between points, and clearly organised arguments. Low-quality responses lack clarity, exhibit abrupt transitions, or are disorganised (Paltridge, 2004). | 0.75 |
| Linguistic Quality | Reflects grammar accuracy, vocabulary appropriateness, sentence fluency, and overall clarity of written expression. High-quality responses utilize clear, precise, and varied language with minimal errors, while low-quality responses contain grammatical errors, limited vocabulary, or unclear expression (Biber and Gray, 2010). | 0.89 |
| Critical Thinking | Evaluates the level of careful thought, complexity, consideration of alternative viewpoints, and strength of conclusions and recommendations. High-quality responses deeply analyze complexities and implications, considering broader contexts, contingencies, or solutions. Low-quality responses show limited or superficial analytical engagement and fail to address alternative perspectives or implications (Facione, 2000; Moon, 2007). | 0.80 |
@@ 3.4.4 Scoring Procedures and Inter-Rater Reliability
The scoring process followed detailed guidelines based on the rubric provided in Appendix C. Two independent, trained coders rated each participant’s writing samples at each study phase (AI-support and AI-removal). Coders were trained extensively on the rubric definitions through practice scoring sessions with sample responses (n=10), followed by discussion and calibration of understanding to maximise rating consistency. All essays were then independently blind-rated by both coders. After ratings were fully completed, reliability analyses were performed (Spearman’s correlations), demonstrating high inter-rater reliability ( in all cases; Table 1). Final participant scores for each dimension, at each stage, were averaged across the two coders, producing a single composite rating for each dimension and subsequently an overall Composite Score. To ensure transparency and replicability, example excerpts representing each scoring level were provided in the detailed rubric (see Appendix C).
@@ 3.5 Study Procedure
We conducted a two-phase, mixed-method experimental study to examine how GenAI literacy and AI interaction type (reactive vs. proactive chatbot) influenced participants’ academic writing performance with visual information. Participants were recruited via Prolific and directed to a Qualtrics222https://www.qualtrics.com/ survey, where they provided informed consent, demographics, and completed literacy assessments (GLAT and mini-VLAT). After reviewing contextual information on the healthcare simulation (Section 3.3.1), they were randomly assigned to either the reactive or proactive chatbot assistance condition. In the AI-supported phase, participants composed a 150-word written response interpreting the provided multimodal visuals with real-time chatbot support. Subsequently, in the AI-removal phase, participants completed a similar task with new visual data from a different simulation scenario without chatbot access (Figure 4). The study lasted approximately one hour. Ethical approval was obtained from [Anonymised] University (Project Number: Anonymised).
@@ 3.6.1 Correlation Analysis (RQ1)
To address RQ1, we conducted a series of correlation analyses to investigate the relationships between students’ GenAI literacy and their multimodal academic writing performance. The analysis proceeded in two stages.
First, we examined the direct associations between literacy scores and writing outcomes. We explored the relationships of both GenAI literacy (as measured by the GLAT) and visualisation literacy (as measured by the mini-VLAT), which served as a proxy for the specific domain knowledge required for the task, with writing performance scores. Prior to analysis, Shapiro-Wilk tests were conducted on all variables, which indicated that the literacy and writing scores were not normally distributed (). Consequently, we employed Spearman’s rank-order correlation () for all subsequent analyses. These correlations were computed separately for the reactive and proactive chatbot conditions at both the AI-support and AI-removal time points. The analysis was applied to the composite writing score as well as to each individual rubric dimension, including insightfulness, visual data integration, organisation and coherence, linguistic quality, and critical thinking.
To provide a more rigorous test of our research question, the second stage of our analysis aimed to isolate the unique contribution of GenAI literacy. We performed a partial correlation analysis to isolate the influence of GenAI literacy from that of task-specific domain knowledge (i.e., visualisation literacy), thereby controlling for potential confounding effects. This involved first computing the zero-order Spearman’s correlation between GLAT and VLAT scores to assess their degree of overlap within each condition. Subsequently, we conducted a semi-partial correlation to determine the association between GenAI literacy and writing performance after statistically removing the variance attributable to visualisation literacy. This conservative approach allowed for a more precise estimation of the unique relationship between GenAI literacy and writing performance. This analysis was conducted consistently with the primary correlations, examining each chatbot condition and time point separately. For all analyses, statistical significance was set at .
@@ 3.6.2 Comparative Analysis (RQ2)
To investigate whether students with varying levels of GenAI literacy performed differently in academic writing (RQ2), we conducted a series of comparative analyses. As a preliminary step, we first examined whether domain knowledge was a significant confounding variable in relation to writing performance. To do this, we examined if differences in visualisation literacy were associated with writing performance. Participants in each condition were divided into low and high visualisation literacy groups based on a median split of their VLAT scores. We chose the median split procedure to ensure balanced group sizes and improve interpretability. Mann-Whitney tests were then performed to compare the writing scores between these groups, which are appropriate for the ordinal nature of the rubric scales and the non-normal distribution of the data (verified through Shapiro-Wilk tests). This initial analysis allowed us to determine whether domain knowledge was a significant factor in its own right before proceeding to our primary question.
Having established the role of domain knowledge, we conducted the main analysis for RQ2, to investigate if students with varying level of GenAI literacy performed differently in academic writing. Participants were again categorised into low and high literacy groups, this time based on a median split of their GLAT scores within each condition. Mann-Whitney tests were also selected for this analysis.
For both analyses of the VLAT and GLAT scores, tests were conducted separately for each rubric dimension (insightfulness, visual data integration, organisation and coherence, linguistic quality, and critical thinking) as well as for the composite score, across both the AI-support and AI-removal stages. Statistical significance was set at and adjusted for multiple comparisons using the Bonferroni method. Significant differences were further interpreted using effect sizes calculated as the Rank-Biserial correlation.
@@ 3.6.3 Regression Analysis (RQ3)
To investigate the extent GenAI literacy predicted writing performance after controlling for domain-specific knowledge (RQ3), ordinal logistic regression analyses were conducted. As shown in the following formula (2), the dependent variables for regression analyses were participants’ academic writing performance scores during the AI-removal period for each dimension (insightfulness, visual data integration, organisation and coherence, linguistic quality, critical thinking) separately, as well as the composite writing score. Each outcome variable, measured using ordinal scales (; scores from 1 to 5; refers to each ordinal threshold, ), justified the use of ordinal logistic regression rather than linear regression. Primary predictor variables included standardised GenAI literacy scores (), domain knowledge (; VLAT score), chatbot intervention type (; binary-coded: 0 = reactive chatbot, 1 = proactive chatbot), and participants’ AI-assisted scores (; to control writing status during chatbot support availability). Additionally, denotes the threshold/intercept parameter for each category boundary, and , and represent regression coefficients for the predictors.
@@
3.6.4 Mediation Analysis (RQ4)
To investigate the mechanism through which GenAI literacy influences independent writing performance (RQ4), we conducted a mediation analysis. The primary hypothesis was that the relationship between a student’s GenAI literacy (the independent variable, X) and their subsequent independent writing performance at the AI-removal stage (the outcome variable, Y) would be mediated by their writing performance during the AI-support stage (the mediator, M). This analysis was designed to test whether GenAI literacy enhances in-task performance, which in turn transfers to improved independent writing capabilities. The mediation analysis was specified using two linear models. The mediation model (Eq. 3) estimated the effect of GenAI literacy on AI-support performance, while the outcome model (Eq. 4) estimated the effect of both GenAI literacy and AI-support performance on independent writing performance. Here, captures the effect of GenAI literacy on AI-support performance, captures the effect of AI-support performance on independent writing performance controlling for GenAI literacy, and represents the direct effect of GenAI literacy on independent writing performance. The indirect effect is given by .
We employed a modern bootstrapping procedure with 5,000 resamples to estimate the indirect effect () and its 95% confidence interval. This method was selected for its statistical power and because it does not assume a normal sampling distribution for the indirect effect, making it more robust than traditional tests. The analysis was performed separately for the reactive and proactive chatbot conditions to examine whether the mediational pathway differed based on the interaction strategy. A significant indirect effect was determined if the 95% confidence interval did not contain zero.